# 25mm25mm25mm26mm

### v2

@@ 1. Introduction
Despite the central role of the overlaps, rigorous results about their size and distribution have appeared only recently. Inspired by the seminal work of Chalker and Mehlig [21] computing the expectations of the $\mathcal{O}_{ij}$ for the complex Ginibre ensemble, Bourgade and Dubach [16] also computed the second moment of the off-diagonal overlap,
@@
Is this quadratic decay ubiquitous, in particular does it extend to general real and complex i.i.d. matrix ensembles without invariance properties? In this paper we answer affirmatively by proving that the bound
@@
holds uniformly in the spectrum with very high probability for any small $\xi>0$. The optimality (modulo the $n^{\xi}$ factor) of (1.4) follows, for example, from (1.3). Previously, only two non-optimal bounds were available. In [28, Theorem 3.1] it was shown that $|\mathcal{O}_{ij}|/\sqrt{\mathcal{O}_{ii}\mathcal{O}_{jj}}\leq n^{-\epsilon}$ for bulk eigenvectors, if $|\sigma_{i}-\sigma_{j}|\geq n^{-1/2+\epsilon\textquoteright}$ which is off by almost an entire factor $n$ for far away eigenvalues. Close to the edge of the spectrum, $|\sigma_{i}|,|\sigma_{j}|\approx 1$, half of the optimal decay, $1/(\sqrt{n}|\sigma_{i}-\sigma_{j}|)$, was obtained in [31, Corollary 3.4]. Neither proof can be directly improved and the optimal bound (1.4) requires to resolve two new challenges.
We now explain the novelties in our proof of (1.4). The main inputs are the so-called multi–resolvent local laws for products of resolvents of the Hermitisation of $X-z$. Since the non-Hermitian eigenvectors $\mathbf{l}_{i},\mathbf{r}_{i}$ are not accessible directly, we relate them to the eigenvectors $\mathbf{w}_{i}^{z}=((\mathbf{u}_{i}^{z})^{*},(\mathbf{v}_{i}^{z})^{*})^{*}\in{\mathbb{C}}^{2n}$ of the $(2n)\times(2n)$ Hermitized matrix
@@
Note that $\mathbf{u}_{i}^{z},\mathbf{v}_{i}^{z}\in{\mathbb{C}}^{n}$ are also the left/right singular vectors corresponding to the singular value $\lambda_{i}^{z}$ of a family of matrices $X-z$ parametrized by $z\in{\mathbb{C}}$. We assume that the singular values $\lambda_{i}^{z}$ are labelled in an increasing order.
Eigenvectors of $X$ and singular vectors of $X-z$ are not related in general, except in the special case when $z$ is an eigenvalue. Simple algebra shows that
@@
and in this case the eigenvectors of $X$ with eigenvalue $z$ and the singular vectors of $X-z$ with singular value zero coincide. This fundamental relation will be our starting point to study the eigenvalues and the eigenvectors of the non–Hermitian $X$ via the study of the spectrum of the Hermitized matrix $H^{z}$. Our goal is thus to estimate quantities of the form $|\langle\mathbf{u}_{1}^{z_{1}},\mathbf{u}_{1}^{z_{2}}\rangle|,|\langle\mathbf{v}_{1}^{z_{1}},\mathbf{v}_{1}^{z_{2}}\rangle|$ with high probability simultaneously for every $|z_{1}|,|z_{2}|\leq 1$. In fact, if we have the desired bound on these quantities, by choosing $z_{1}=\sigma_{i}$, $z_{2}=\sigma_{j}$, we can readily conclude (1.4). By spectral decomposition, one can easily see that for the desired bound on $|\langle\mathbf{u}_{1}^{z_{1}},\mathbf{u}_{1}^{z_{2}}\rangle|,|\langle\mathbf{v}_{1}^{z_{1}},\mathbf{v}_{1}^{z_{2}}\rangle|$, it is actually enough to estimate
@@
We close this introduction with a quick overview of the recent vast literature about overlaps (1.2) for various random matrix ensembles. With the exception of [16], all these works concern only diagonal overlaps $\mathcal{O}_{ii}$. Long after the classical paper [21], the two main works that revived the interest in overlaps of non–Hermitian eigenvectors in the context of random matrix theory are [16, 46] on Ginibre ensemble. Besides the results on the off-diagonal overlaps in [16] that we already described, the distribution of $\mathcal{O}_{ii}$ and the correlation of $\mathcal{O}_{ii}\mathcal{O}_{jj}$ for bulk eigenvalues in the complex case have also been computed in that paper. In [46] the author computes the distribution of $\mathcal{O}_{ii}$ both in the bulk and at the edge of the spectrum for complex matrices, as well as for overlaps corresponding to real eigenvalues of real Ginibre matrices. Expectation of overlaps for complex eigenvalues was covered in [71]. Beyond Gaussian ensembles, in the context of general i.i.d. matrices, the precise (in terms of $n$–dependence) size of $\mathcal{O}_{ii}$ was identified in [22, 44] (see also for previous non-optimal results in terms of the $n$–dependence [10, 58, 73]), and more recently even the universality of its distribution has been established [67]. Two very recent results [43, 66] also prove the universality of the Gaussian fluctuations directly for the eigenvectors’ entries of i.i.d. matrices. Eigenvector overlaps have also been studied for several other models even beyond i.i.d. matrices, including deformed Ginibre matrices [11, 69, 75], symplectic Ginibre matrices [5, 4], non–Hermitian Brownian motion [15, 45], the elliptic ensemble [37, 38, 51, 72], spherical and truncated unitary ensembles [40, 41, 64], and non-normal dynamical systems [63].
@@ Acknowledgment
We thank Yan Fyodorov for pointing out the reference [52] and for useful comments.
@@
For integers $k,l\in{\mathbb{N}}$, with $k<l$, we use the notations $[k,l]:=\{k,\dots,l\}$ and $[k]:=[1,k]$. For positive quantities $f,g$ we write $f\lesssim g$ and $f\sim g$ if $f\leq Cg$ or $cg\leq f\leq Cg$, respectively, for some constants $c,C>0$ which depend only on the constants appearing in (2.1). We set $f\wedge g=\min\{f,g\}$ and $f\vee g=\max\{f,g\}$. Furthermore, for $n$–dependent positive quantities $f_{n},g_{n}$ we use the notation $f_{n}\ll g_{n}$ to denote that $\lim_{n\to\infty}(f_{n}/g_{n})=0$. Throughout the paper $c,C>0$ denote small and large constants, respectively, which may change from line to line. We denote vectors by bold-faced lower case Roman letters $\bm{x},\bm{y}\in{\mathbb{C}}^{d}$, for some $d\in{\mathbb{N}}$. Vector and matrix norms, $\lVert\bm{x}\rVert$ and $\lVert A\rVert$, indicate the usual Euclidean norm and the corresponding induced matrix norm. For any $d\times d$ matrix $A$ we use $\langle A\rangle:=d^{-1}\mathrm{Tr}A$ to denote the normalized trace of $A$. Moreover, for vectors ${\bm{x}},{\bm{y}}\in{\mathbb{C}}^{d}$ we define the usual scalar product $\langle\bm{x},\bm{y}\rangle:=\sum_{i=1}^{d}\overline{x_{i}}y_{i}$, and use $(A)_{{\bm{x}}{\bm{y}}}$ to denote the inner product $\langle{\bm{x}},A{\bm{y}}\rangle$.
@@
in particular, $E_{+}$ is the $2n$–dimensional identity matrix.
@@ 2. Main results
Consider an i.i.d. matrix $X$ satisfying Assumption 2.1. Furthermore, we assume that $\chi$ has a bounded density $g$ and in the complex case $\mathrm{Re}\,\chi,\mathrm{Im}\,\chi$ are independent. Fix any small constants $\tau,\epsilon,\xi>0$ with $\xi\leq\epsilon/10$ and any large constant $D>0$. Then for any $z_{1},z_{2}\in{\mathbb{C}}$ with with $|z_{1}-z_{2}|\geq n^{-1/2+\epsilon}$, there is an event $\Xi$ with very high probability, i.e. ${\mathbf{P}}(\Xi)\geq 1-n^{-D}$, such that
[…]
hold for any indices $i,j$ with the corresponding eigenvalues close to $z_{1}$ and $z_{2}$, respectively, i.e. such that $|\sigma_{i}-z_{1}|+|\sigma_{j}-z_{2}|\leq n^{-1/2+\xi}$.
Furthermore, in the complex case in the bulk regime, i.e if $|z_{1}|,|z_{2}|\leq 1-\tau$, the event $\Xi$ is trivial, i.e. (2.6) holds without ${\bf 1}_{\Xi}$. In the real case we have a special bound for eigenvalues $\sigma_{i},\sigma_{j}$ near $z_{1},z_{2}$ on the real axis:
${\mathbf{E}}\Big{(}|\mathcal{O}_{ij}|^{1/2}\Big{)}\lesssim\frac{n^{\xi}}{|z_{1}-z_{2}|},\qquad\sigma_{i},\sigma_{j}\in{\mathbb{R}},\quad|\sigma_{i}-z_{1}|+|\sigma_{j}-z_{2}|\leq n^{-1/2+\xi}.$ (2.8)
@@ 3.1. Local laws for the resolvent
Note that $z\in{\mathbb{C}}$ lies in the spectrum of $X$ if only and if the least singular value of $X-z$ is zero. Hence we link the eigenvalues (resp. eigenvectors) of $X$ with the singular values (resp. singular vectors) of matrix $X-z$. As already mentioned in the Introduction, let $(\lambda^{z}_{i})_{i=1}^{n}$ denote the singular values of $X-z$ labelled in a non-decreasing order, and $\{\mathbf{u}_{i}^{z}\}_{i=1}^{n},\{\mathbf{v}_{i}^{z}\}_{i=1}^{n}$ denote the corresponding left and right singular vectors of $X-z$ normalised such that $\|\mathbf{u}^{z}_{i}\|^{2}=\|\mathbf{v}^{z}_{i}\|^{2}=1/2$. In the same spirit of [54], we consider the Hermitisation of $X$, i.e.,
@@
Simple algebra (also called chiral symmetry) shows that $H^{z}$ has a symmetric spectrum given by $(\lambda^{z}_{\pm i})_{i=1}^{n}=(\pm\lambda^{z}_{i})_{i=1}^{n}$, with the corresponding (orthonormal) eigenvectors given by $\mathbf{w}^{z}_{\pm i}=((\mathbf{u}^{z}_{i})^{*},\pm(\mathbf{v}^{z}_{i})^{*})^{*}\in{\mathbb{C}}^{2n}$. Note that $\mathbf{u}^{z}_{i},\mathbf{v}^{z}_{i}$ are the left and right singular vectors of $X-z$. Eigenvalues of $H^{z}$ in general have no direct relation with the eigenvalues of $X$, however, from the trivial relation (1.6) it is clear that eigenvectors of $X$ with eigenvalue $z$ and the singular vectors of $H^{z}$ with singular value zero coincide.
Let $G^{z}$ be the resolvent of the Hermitisation of $X-z$, i.e.,
@@
$-[M^{z}(w)]^{-1}=w+Z+\mathcal{S}[M^{z}(w)],\qquad Z:=\left(\begin{matrix}0&z\\ \overline{z}&0\end{matrix}\right),$ (3.5)
@@
    if $|z|>1$, then there exists a small gap $[-\Delta/2,\Delta/2]$ in the support of the symmetric density $\rho^{z}(E)$, with $\Delta\sim(|z|^{2}-1)^{3/2}$.
@@
|#$\Big{|}\#\{j:E_{1}\leq\lambda^{z}_{j}\leq E_{2}\}-2n\int_{E_{1}}^{E_{2}}\rho^{z}(x)\mathrm{d}x\Big{|}\prec 1,\qquad|E_{i}|\leq c^{\prime},\quad i=1,2.$ (3.15)

### v3

# Optimal decay of eigenvector overlap for non-Hermitian random matrices

@@ 1. Introduction
Despite the central role of the overlaps, rigorous results about their size and distribution have appeared only recently. Inspired by the seminal work of Chalker and Mehlig [21] computing the expectations of the $\mathcal{O}_{ij}$ for the complex Ginibre ensemble, Bourgade and Dubach [16] also computed the second moment of the off-diagonal overlap and diagonal overlap,
@@
${\mathbf{E}}\big[\mathcal{O}_{ii}\mathcal{O}_{jj}\big|\sigma_{i}=z_{1},\,\,\sigma_{j}=z_{2}\big]\approx n^{2}(1-|z_{1}|^{2})(1-|z_{2}|^{2}),$ (1.4)
@@
Is this quadratic decay ubiquitous, in particular does it extend to general real and complex i.i.d. matrix ensembles without invariance properties? In this paper we answer affirmatively by proving that the optimal bound (modulo the $n^{\xi}$ factor) as indicated by (1.3)-(1.4)
@@
holds uniformly in the spectrum with very high probability for any small $\xi>0$. Previously, only two non-optimal bounds were available. In [28, Theorem 3.1] it was shown that $|\mathcal{O}_{ij}|/\sqrt{\mathcal{O}_{ii}\mathcal{O}_{jj}}\leq n^{-\epsilon}$ for bulk eigenvectors, if $|\sigma_{i}-\sigma_{j}|\geq n^{-1/2+\epsilon’}$ which is off by almost an entire factor $n$ for far away eigenvalues. Close to the edge of the spectrum, $|\sigma_{i}|,|\sigma_{j}|\approx 1$, half of the optimal decay, $1/(\sqrt{n}|\sigma_{i}-\sigma_{j}|)$, was obtained in [31, Corollary 3.4]. We remark that the quadratic decay bound obtained in (1.5) is better than $1/(\sqrt{n}|\sigma_{i}-\sigma_{j}|)$ if $|\sigma_{i}-\sigma_{j}|\geq n^{-1/2+\xi}$. Neither proof can be directly improved and the optimal bound (1.5) requires to resolve two new challenges.
We now explain the novelties in our proof of (1.5). The main inputs are the so-called multi–resolvent local laws for products of resolvents of the Hermitisation of XzX-z. Since the non-Hermitian eigenvectors $\mathbf{l}_{i},\mathbf{r}_{i}$ are not accessible directly, we relate them to the eigenvectors $\mathbf{w}_{i}^{z}=((\mathbf{u}_{i}^{z})^{*},(\mathbf{v}_{i}^{z})^{*})^{*}\in{\mathbb{C}}^{2n}$ of the $(2n)\times(2n)$ Hermitized matrix
@@
Note that $\mathbf{u}_{i}^{z},\mathbf{v}_{i}^{z}\in{\mathbb{C}}^{n}$ are also the left/right singular vectors corresponding to the singular value $\lambda_{i}^{z}$ of a family of matrices XzX-z parametrized by $z\in{\mathbb{C}}$. We assume that the singular values $\lambda_{i}^{z}$ are labelled in an increasing order.
Eigenvectors of $X$ and singular vectors of XzX-z are not related in general, except in the special case when $z$ is an eigenvalue. Simple algebra shows that
@@
and in this case the eigenvectors of $X$ with eigenvalue $z$ and the singular vectors of XzX-z with singular value zero coincide. This fundamental relation will be our starting point to study the eigenvalues and the eigenvectors of the non–Hermitian $X$ via the study of the spectrum of the Hermitized matrix $H^{z}$. Our goal is thus to estimate quantities of the form $|\langle\mathbf{u}_{1}^{z_{1}},\mathbf{u}_{1}^{z_{2}}\rangle|,|\langle\mathbf{v}_{1}^{z_{1}},\mathbf{v}_{1}^{z_{2}}\rangle|$ with high probability simultaneously for every $|z_{1}|,|z_{2}|\leq 1$. In fact, if we have the desired bound on these quantities, by choosing $z_{1}=\sigma_{i}$, $z_{2}=\sigma_{j}$, we can readily conclude (1.5). By spectral decomposition, one can easily see that for the desired bound on $|\langle\mathbf{u}_{1}^{z_{1}},\mathbf{u}_{1}^{z_{2}}\rangle|,|\langle\mathbf{v}_{1}^{z_{1}},\mathbf{v}_{1}^{z_{2}}\rangle|$, it is actually enough to estimate
@@
We close this introduction with a quick overview of the recent vast literature about overlaps (1.2) for various random matrix ensembles. With the exception of [16], all these works concern only diagonal overlaps $\mathcal{O}_{ii}$. Long after the classical paper [21], the two main works that revived the interest in overlaps of non–Hermitian eigenvectors in the context of random matrix theory are [16, 46] on Ginibre ensemble. Besides the results on the off-diagonal overlaps in [16] that we already described, the distribution of $\mathcal{O}_{ii}$ and the correlation of $\mathcal{O}_{ii}\mathcal{O}_{jj}$ for bulk eigenvalues in the complex case have also been computed in that paper. In [46] the author computes the distribution of $\mathcal{O}_{ii}$ both in the bulk and at the edge of the spectrum for complex matrices, as well as for overlaps corresponding to real eigenvalues of real Ginibre matrices. Expectation of overlaps for complex eigenvalues was covered in [71]. Beyond Gaussian ensembles, in the context of general i.i.d. matrices, the precise (in terms of nn–dependence) size of $\mathcal{O}_{ii}$ was identified in [22, 44] (see also for previous non-optimal results in terms of the nn–dependence [10, 58, 73]), and more recently even the universality of its distribution has been established [67]. Two very recent results [43, 66] also prove the universality of the Gaussian fluctuations directly for the eigenvectors’ entries of i.i.d. matrices. Eigenvector overlaps have also been studied for several other models even beyond i.i.d. matrices, including deformed Ginibre matrices [11, 69, 76], symplectic Ginibre matrices [5, 4], non–Hermitian Brownian motion [15, 45], the elliptic ensemble [37, 38, 51, 72], spherical and truncated unitary ensembles [40, 41, 64], and non-normal dynamical systems [63].
@@ Acknowledgment
We thank Yan Fyodorov for pointing out the reference [52] and for useful comments. We also thank the anonymous referee for their several comments that substantially improved the presentation of this manuscript.
@@
For integers $k,l\in{\mathbb{N}}$, with $k<l$, we use the notations $k,l]:=\{k,\dots,l\}$ and $[k]:=[1,k]$. For positive quantities $f,g$ we write $f\lesssim g$ and $f\sim g$ if $f\leq Cg$ or $cg\leq f\leq Cg$, respectively, for some constants $c,C>0$ which depend only on the constants appearing in (2.1). We set $f\wedge g=\min\{f,g\}$ and $f\vee g=\max\{f,g\}$. Furthermore, for nn–dependent positive quantities $f_{n},g_{n}$ we use the notation $f_{n}\ll g_{n}$ to denote that $\lim_{n\to\infty}(f_{n}/g_{n})=0$. Throughout the paper $c,C>0$ denote small and large constants, respectively, which may change from line to line. We denote vectors by bold-faced lower case Roman letters $\bm{x},\bm{y}\in{\mathbb{C}}^{d}$, for some $d\in{\mathbb{N}}$. Vector and matrix norms, $\lVert\bm{x}\rVert$ and $\lVert A\rVert$, indicate the usual Euclidean norm and the corresponding induced matrix norm. For any $d\times d$ matrix $A$ we use $\langle A\rangle:=d^{-1}\mathrm{Tr}A$ to denote the normalized trace of $A$. Moreover, for vectors ${\bm{x}},{\bm{y}}\in{\mathbb{C}}^{d}$ we define the usual scalar product $\langle\bm{x},\bm{y}\rangle:=\sum_{i=1}^{d}\overline{x_{i}}y_{i}$, and use $(A)_{{\bm{x}}{\bm{y}}}$ to denote the inner product $\langle{\bm{x}},A{\bm{y}}\rangle$.
@@
in particular, $E_{+}$ is the 2n2n–dimensional identity matrix.
@@ 2. Main results
Remark 2.4.
In Theorem 2.2 we proved an (almost) optimal bound for the overlaps of matrices $X$ with i.i.d. entries as in Assumption 2.1. However, inspecting our proof it is clear that this argument can easily be extended to more general models, i.e. to deformed models $X+A$, where $X$ is an i.i.d. matrix and $A$ is a deterministic deformation belonging to a large class of matrices (see e.g. [22, Section 2.1], [20, Assumption 2.8] for bulk and edge eigenvectors, respectively). More precisely, the arguments of Sections 45 work verbatim in this more general case as well, once the deterministic bounds in Lemma 3.3 are established.
@@
Consider an i.i.d. matrix $X$ satisfying Assumption 2.1. Furthermore, we assume that $\chi$ has a bounded density $g$ and in the complex case $\mathrm{Re}\,\chi,\mathrm{Im}\,\chi$ are independent. Fix any small constants $\tau,\epsilon,\xi>0$ with $\xi\leq\epsilon/10$ and any large constant $D>0$. Set $\mathcal{D}^{\xi}_{z}\subset{\mathbb{C}}$ to be the disk centered at $z\in{\mathbb{C}}$ with radius $n^{-1/2+\xi}$. Then for any $z_{i}\in{\mathbb{C}}~(i=1,2)$ with $|z_{i}|\leq 1+n^{-1/2+\tau}$ and $|z_{1}-z_{2}|\geq n^{-1/2+\epsilon}$, there is an event $\Xi$ with very high probability, i.e. ${\mathbf{P}}(\Xi)\geq 1-n^{-D}$, such that
${\mathbf{E}}\Big(\sum_{\sigma_{i}\in\mathcal{D}^{\xi}_{z_{1}},\sigma_{j}\in\mathcal{D}^{\xi}_{z_{2}}}|\mathcal{O}_{ij}|\cdot{\bf 1}_{\Xi}\Big)$ $\lesssim\frac{n^{8\xi}}{|z_{1}-z_{2}|^{2}},\qquad\mbox{[Complex case]}$ (2.6)
${\mathbf{E}}\Big(\sum_{\sigma_{i}\in\mathcal{D}^{\xi}_{z_{1}},\sigma_{j}\in\mathcal{D}^{\xi}_{z_{2}}}|\mathcal{O}_{ij}|\cdot{\bf 1}_{\Xi}\Big)$ $\lesssim\frac{n^{8\xi}}{|z_{1}-z_{2}|^{2}}\cdot\frac{1}{|\mathrm{Im}\,z_{1}|\wedge|\mathrm{Im}\,z_{2}|},\qquad\mbox{[Real case]}$ (2.7)
Furthermore, in the complex case in the bulk regime, i.e if $|z_{1}|,|z_{2}|\leq 1-\tau$, the event $\Xi$ is trivial, i.e. (2.6) holds without ${\bf 1}_{\Xi}$. In the real case we have a special bound for the real eigenvalues $\sigma_{i},\sigma_{j}$ near $z_{1},z_{2}$:
${\mathbf{E}}\Big(\sum_{\begin{subarray}{c}\sigma_{i}\in\mathcal{D}^{\xi}_{z_{1}}\cap{\mathbb{R}}\\ \sigma_{j}\in\mathcal{D}^{\xi}_{z_{2}}\cap{\mathbb{R}}\end{subarray}}|\mathcal{O}_{ij}|^{1/2}\Big)\lesssim\frac{n^{4\xi}}{|z_{1}-z_{2}|}.$ (2.8)
Remark 2.6.
From the local circular laws (see [17, Theorem 2.2], [18, Theorem 1.3], [74, Theorem 1.2], and [9, Theorem 2.3]), we know that with very high probability there are eigenvalues close to $z_{1}$ and $z_{2}$. More precisely, fix any small $\tau,\xi>0$ with $\tau<\xi/10$, then for any $|z_{i}|\leq 1+n^{-1/2+\tau}$, there exist eigenvalues $\sigma_{i}\in\mathcal{D}^{\xi}_{z_{1}}$ and $\sigma_{j}\in\mathcal{D}^{\xi}_{z_{2}}$ with very high probability. Hence, the summations in (2.6)–(2.7) are not empty, in fact they typically contain roughly $n^{2\xi}\times n^{2\xi}$ eigenvalue pairs. Our result holds for the corresponding overlap for any of these pairs.
@@ 3.1. Local laws for the resolvent
Note that $z\in{\mathbb{C}}$ lies in the spectrum of $X$ if only and if the least singular value of XzX-z is zero. Hence we link the eigenvalues (resp. eigenvectors) of $X$ with the singular values (resp. singular vectors) of matrix XzX-z. As already mentioned in the Introduction, let $(\lambda^{z}_{i})_{i=1}^{n}$ denote the singular values of XzX-z labelled in a non-decreasing order, and $\{\mathbf{u}_{i}^{z}\}_{i=1}^{n},\{\mathbf{v}_{i}^{z}\}_{i=1}^{n}$ denote the corresponding left and right singular vectors of XzX-z normalised such that $\|\mathbf{u}^{z}_{i}\|^{2}=\|\mathbf{v}^{z}_{i}\|^{2}=1/2$. In the same spirit of [54], we consider the Hermitisation of $X$, i.e.,
@@
Simple algebra (also called chiral symmetry) shows that $H^{z}$ has a symmetric spectrum given by $(\lambda^{z}_{\pm i})_{i=1}^{n}=(\pm\lambda^{z}_{i})_{i=1}^{n}$, with the corresponding (orthonormal) eigenvectors given by $\mathbf{w}^{z}_{\pm i}=((\mathbf{u}^{z}_{i})^{*},\pm(\mathbf{v}^{z}_{i})^{*})^{*}\in{\mathbb{C}}^{2n}$. Note that $\mathbf{u}^{z}_{i},\mathbf{v}^{z}_{i}$ are the left and right singular vectors of XzX-z. Eigenvalues of $H^{z}$ in general have no direct relation with the eigenvalues of $X$, however, from the trivial relation (1.7) it is clear that eigenvectors of $X$ with eigenvalue $z$ and the singular vectors of $H^{z}$ with singular value zero coincide.
Let $G^{z}$ be the resolvent of the Hermitisation of XzX-z, i.e.,
@@
$M^{z}(w)]^{-1}=w+Z+\mathcal{S}[M^{z}(w)],\qquad Z:=\left(\begin{matrix}0&z\\ \overline{z}&0\end{matrix}\right),$ (3.5)
@@
    if $|z|>1$, then there exists a small gap $\Delta/2,\Delta/2]$ in the support of the symmetric density $\rho^{z}(E)$, with $\Delta\sim(|z|^{2}-1)^{3/2}$.
@@
$\Big|\#\{j:E_{1}\leq\lambda^{z}_{j}\leq E_{2}\}-2n\int_{E_{1}}^{E_{2}}\rho^{z}(x)\mathrm{d}x\Big|\prec 1,\qquad|E_{i}|\leq c^{\prime},\quad i=1,2.$ (3.15)